# Resonance Contributions to Charged-Current (Anti)Neutrino-Nucleon Elastic Scattering at GeV Energies

### v1

@@
Resonance Contributions to Charged-Current (Anti)Neutrino-Nucleon Elastic Scattering at GeV Energies
@@
(September 28, 2025)
We present the first evaluation of virtual resonance contributions to the charged-current (anti)neutrino-nucleon elastic scattering at GeV energies, focusing on the dominant $\Delta(1232)$ resonance. We approximate the vector part of the $N\to\Delta$ transition by the leading magnetic dipole term. Our results for the cross-section corrections at fixed neutrino energy indicate the permille-level contribution of resonance intermediate states to the elastic and radiative scattering cross sections. This calculation exhibits the expected infrared behavior of the invariant amplitudes and unpolarized cross sections. Our findings provide important insights into inelastic excitations in the charged-current (anti)neutrino-nucleon elastic scattering at GeV energies.
@@ 1 Introduction
Unpolarized elastic electron scattering is a well-established tool for studying electromagnetic properties of nucleons and nuclei [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Modern measurements of elastic electron-proton scattering at GeV energies and below are performed with precision at the percent level or better. Such accuracy calls for a careful treatment of radiative corrections [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] and nucleon excitations. In particular, the diagram with two exchanged photons is a subject of active theoretical and experimental investigation [25, 26, 27, 28]. It has attracted increasing attention [25, 29, 30] after the first extraction of the proton electromagnetic form-factor ratio from polarization transfer data [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], which resulted in contradiction with a traditional Rosenbluth separation technique at GeV electron beam energies. At these energies, on top of effective field theory considerations [42, 43, 44, 45, 46, 47, 48, 49, 50], there are two well-developed ways to approach the two-photon exchange diagram: hadronic models and dispersion relations. Calculations within the hadronic model [29, 51, 52, 53, 42, 54, 55, 56, 57] assume the on-shell form for the photon-proton interaction and perform a straightforward evaluation of the loop diagram. Such an evaluation is ultraviolet-finite and collinear-safe, but it violates unitarity for the real parts of the two-photon exchange amplitudes. This violation is not significant for the proton, or elastic, intermediate state, when the lepton mass is neglected [58, 59], but it leads to cross sections that diverge with beam energy when modeling the resonance intermediate states [60, 61]. The dispersion-relation approach [62, 63, 64, 58, 65, 59], based on the on-shell information only, does not suffer from such pathological behavior and is considered a physically-motivated and favorable framework for the evaluation of the two-photon exchange diagram in electron-proton scattering community [66]. As of today, proton and pion-nucleon intermediate states in the two-photon exchange diagram are included in the dispersion-relation approach. Such a calculation takes experimentally measured proton form factors [4, 5] and invariant amplitudes of pion electroproduction [67, 68] as an input. Accounting for the pion-nucleon intermediate states improves agreement between theoretical predictions and recent experimental data from electron- and positron-proton scattering experiments [69, 70, 71, 72, 73, 74, 75]. The remaining discrepancy might be explained by the resonant enhancement of nucleon excitations at GeV beam energies [76], when the propagator of the resonance intermediate state moves close to its pole position.
@@
A similar diagram with the exchange of two bosons contributes to the (anti)neutrino-nucleon elastic scattering. To fully benefit from experimental advances in the fields of neutrino scattering and oscillation experiments [77, 78, 79], this diagram was recently revisited by a few independent groups [80, 81, 82, 83, 84]. As a result, it was included in a new and complete factorization framework for radiative corrections to the charged-current (anti)neutrino-nucleon elastic scattering in Refs. [83, 84]. The latter references evaluated the nucleon intermediate state and provided a model consistent with the expected soft and collinear behavior. The authors estimated the uncertainty from inelastic intermediate states as a contribution from the diagrams with electromagnetic coupling to the neutron. As in electron scattering, resonance intermediate states in the charged-current (anti)neutrino elastic scattering might be enhanced for the incoming beam energies that move the resonance propagators close to their pole positions. That is why resonance contributions may be relevant for precision analyses of (anti)neutrino scattering data.
@@
In this paper, we perform the first evaluation of the $\Delta(1232)$-resonance contribution to the unpolarized charged-current (anti)neutrino-nucleon elastic scattering at GeV energies. We verify that fixed-$t$ dispersion relations do not apply to resonance contributions as they are for the elastic intermediate state [83, 84]. Therefore, we calculate the loop integrals by modeling the $\Delta$ interaction vertex in its on-shell form with two choices: (1) shifting momenta to satisfy collinear constraints as in the model of Refs. [83, 84] and (2) retaining the full momentum dependence in the loop integration as in the hadronic model of Refs. [29, 51, 52, 53, 54, 55, 56, 57], cf. also an interesting discussion in Ref. [85]. Additionally, we investigate soft and collinear properties of the invariant amplitudes and study off-shell effects by changing the vector part of the weak interaction vertex of the $\Delta$ resonance from the gauge-independent form of the magnetic dipole contribution only to a traditional in neutrino physics form-factor decomposition. Our model satisfies soft and collinear constraints on the real parts of the invariant amplitudes. It also respects the expected crossing properties for several contributions, specified in Section 4, and fulfills the unitarity constraints for the imaginary parts.
@@
The paper is organized as follows. In Section 2, we describe the production of the $\Delta$ resonance in electroweak processes at tree level and specify the amplitudes for the $\Delta$ neutrinoproduction. We provide the invariant amplitude decomposition and general cross-section expression in the (anti)neutrino-nucleon elastic scattering in Section 3. In Section 4, we evaluate the virtual contribution of the $\Delta$ intermediate state to the (anti)neutrino-nucleon elastic scattering amplitudes and cross sections and provide numerical results for incoming (anti)neutrino beam energies around GeV scale. Conclusions and outlook are presented in Section 5. In Appendix A, we specify relations between invariant and helicity amplitudes in the (anti)neutrino-nucleon elastic scattering.
@@ 2 $\Delta$ production in charged-current (anti)neutrino-nucleon scattering
We consider all possible channels for the production of the $\Delta$ resonance in the charged-current scattering of the neutrino $\nu_{\ell}$ and antineutrino $\overline{\nu}_{\ell}$ on the neutron $n$ and proton $p$: 1) $\nu_{\ell}n\to\ell^{-}\Delta^{+}$; 2) $\nu_{\ell}p\to\ell^{-}\Delta^{++}$; 3) $\overline{\nu}_{\ell}n\to\ell^{+}\Delta^{-}$; 4) $\overline{\nu}_{\ell}p\to\ell^{+}\Delta^{0}$. We denote the momentum of the incoming (anti)neutrino as $p$, the momentum of the outgoing charged lepton as $p^{\prime}$, the momentum of the initial-state nucleon as $k$, and the momentum of the final-state resonance as $k^{\prime}$. The kinematics of the resonance production can be completely described by the invariant mass of the resonance $W$, where $W^{2}=\left(k^{\prime}\right)^{2}$ and two Mandelstam variables: the squared momentum transfer $Q^{2}=-\left(k^{\prime}-k\right)^{2}$ and the squared energy in the (anti)neutrino-nucleon center-of-mass reference frame $s=\left(k+p\right)^{2}$.
@@
where the neutron electric form factor is taken from Ref. [109]. For the neutron magnetic, proton electric, and proton magnetic form factors, a dipole form is assumed. Using the magnetic dipole transition approximation in Eq. (5), the coupling to virtual photons becomes gauge-invariant and the unpolarized $\Delta$ production cross section does not contain vector-axial-vector interference terms. In this approximation, the vector transition form factors in Eq. (3) are given by
@@
The conservation of the vector current (CVC) implies $C^{V}_{6}=0$.
@@
which we call the $\Delta_{\mathrm{I}}$ model. We also consider an alternative approximation $C^{A}_{4}\left(Q^{2}\right)=0$ of Ref. [112], which we call the $\Delta_{\mathrm{II}}$ model. For numerical estimates in this paper, we take the axial-vector form factor $C^{A}_{5}\left(Q^{2}\right)$ from Ref. [105]:
@@
3 Charged-current (anti)neutrino-nucleon elastic scattering
In this Section, we describe the invariant amplitude decomposition in the (anti)neutrino-nucleon elastic scattering for the charged-current processes.
@@ 3 Charged-current (anti)neutrino-nucleon elastic scattering
with the averaged nucleon momentum $K_{\mu}=(k_{\mu}+k^{\prime}_{\mu})/2$. Accounting for the nonzero lepton mass, four more invariant amplitudes are required to describe the matrix elements of the charged-current elastic processes, i.e., $T_{\nu_{\ell}n\to\ell^{-}p}=T^{m_{\ell}=0}_{\nu_{\ell}n\to\ell^{-}p}+T^{m_{\ell}\neq 0}_{\nu_{\ell}n\to\ell^{-}p}$ and $T_{\overline{\nu}_{\ell}p\to\ell^{+}n}=T^{m_{\ell}=0}_{\overline{\nu}_{\ell}p\to\ell^{+}n}+T^{m_{\ell}\neq 0}_{\overline{\nu}_{\ell}p\to\ell^{+}n}$:
@@
4 $\Delta$ intermediate state in charged-current (anti)neutrino-nucleon elastic scattering
@@
Figure 1: Contributions of the $\Delta$ resonance to the charged-current (anti)neutrino-nucleon elastic scattering are shown. The photon is exchanged between the charged lepton and $N\to\Delta$ transition vertex.
@@
$\mathfrak{Re}\left[f^{\mathrm{x}}_{A},f^{\mathrm{x}}_{P},f^{\mathrm{x}}_{P5},f^{\mathrm{x}}_{3}\right]\left(\nu,\penalty 10000\ t\right)$ $=\mathfrak{Re}\left[f^{\mathrm{d}}_{A},f^{\mathrm{d}}_{P},f^{\mathrm{d}}_{P5},f^{\mathrm{d}}_{3}\right]\left(-\nu,\penalty 10000\ t\right).$ (30)
@@
The results for $\Delta^{0}_{\mathrm{I}}$ and $\Delta^{0}_{\mathrm{II}}$, calculated using different axial-vector transition form factors, cf. Eq. (12) and the description above are in very close agreement. This indicates the dominant role of the electroweak axial-vector contribution from the form factor $C_{5}^{A}$. The subtraction of power-suppressed collinear contributions has a negligible impact on the results, which validates the approximate flavor independence for the resonance contributions to the unpolarized cross section. For the electron flavor, this subtraction is negligible compared to the magnitude of the $\Delta$-resonance contribution. However, it is more significant for the muon flavor due to power corrections that involve the muon mass. To explicitly study the mass dependence, we subtract the power-suppressed collinear region from each loop integral in the default model while keeping the kinematic relations between the amplitudes and loop integrals exact. We refer to this calculation as $\Delta^{0}_{\mathrm{I}},\penalty 10000\ \mathrm{no\penalty 10000\ coll.}$”, and compare it with the results obtained without any subtraction.
Nevertheless, the $\Delta(1232)$-resonance contribution demonstrates an expected kinematic enhancement, as shown in Figures 2-5, the resulting correction does not exceed the permille level. For the dominant region of (anti)neutrino fluxes in oscillation experiments, the kinematic enhancement happens away from the near-forward region, $Q^{2}\approx 0$, with the largest cross sections. For further validation, we also perform a calculation in the hadronic model $\Delta^{\mathrm{hm}}_{\mathrm{I},\mathrm{II}}$ and find a permille-level correction with a larger kinematic enhancement in a relatively limited phase-space region. The differences between model calculations for the $\Delta$-resonance contribution to the unpolarized (anti)neutrino-nucleon elastic scattering are consistent with the uncertainty estimates in Refs. [83, 84].
@@
Figure 2: Relative contribution of the $\Delta$ intermediate state to the unpolarized charged-current elastic scattering cross section on nucleons is shown as a function of the squared momentum transfer $Q^{2}$ for initial neutrino energies $E_{\nu}=600\penalty 10000\ \mathrm{MeV}$ (upper left plot), $E_{\nu}=1\penalty 10000\ \mathrm{GeV}$ (upper right plot), $E_{\nu}=1.2\penalty 10000\ \mathrm{GeV}$ (lower left plot), and $E_{\nu}=2\penalty 10000\ \mathrm{GeV}$ (lower right plot). Results for different model calculations, which are described in Section 4, are presented.
Figure 3: Same as Figure 2 but for the antineutrino scattering.
@@
Given the permille-level virtual correction from resonance intermediate states, we do not expect larger effects from diagrams involving real photon radiation. Our study also finds a 11-33 permille-level dependence on the treatment of off-shell contributions in the radiative $\Delta$ decay, which rules out predictions based on the decomposition of vector contributions in Eq. (3) more precise than order-of-magnitude estimates for both moderate and large momentum transfers. This result is in qualitative agreement with findings for the radiative $\Delta$ decay in neutral-current process [121, 122]. We therefore provide an independent uncertainty estimate for background channels in searches for new physics at short-baseline neutrino experiments [123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147].

### v3

# Resonance Contributions to Radiative Corrections in Charged-Current Elastic (Anti)Neutrino-Nucleon Scattering at GeV Energies

@@
Resonance Contributions to Radiative Corrections in Charged-Current Elastic (Anti)Neutrino-Nucleon Scattering at GeV Energies
@@
(January 26, 2026)
We present the first evaluation of virtual resonance contributions to the charged-current elastic (anti)neutrino-nucleon scattering at GeV energies, focusing on the dominant $\Delta(1232)$ resonance. We approximate the vector part of the $N\to\Delta$ transition by the leading magnetic dipole term. Our results for the cross-section corrections at fixed neutrino energy indicate the permille-level contribution of resonance intermediate states to the elastic and radiative scattering cross sections. This calculation exhibits the expected infrared behavior of the invariant amplitudes and unpolarized cross sections. Our findings provide important insights into inelastic excitations in the charged-current elastic (anti)neutrino-nucleon scattering at GeV energies.
@@ 1 Introduction
Unpolarized electron elastic scattering is a well-established tool for studying electromagnetic properties of nucleons and nuclei [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Modern measurements of electron-proton elastic scattering at GeV energies and below are performed with precision at the percent level or better. Such accuracy calls for a careful treatment of radiative corrections [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] and nucleon excitations. In particular, the diagram with two exchanged photons is a subject of active theoretical and experimental investigation [25, 26, 27, 28]. It has attracted increasing attention [25, 29, 30] after the first extraction of the proton electromagnetic form-factor ratio from polarization transfer data [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], which resulted in contradiction with a traditional Rosenbluth separation technique at GeV electron beam energies. At these energies, on top of effective field theory considerations [42, 43, 44, 45, 46, 47, 48, 49, 50], there are two well-developed ways to approach the two-photon exchange diagram: hadronic models and dispersion relations. Calculations within the hadronic model [29, 51, 52, 53, 42, 54, 55, 56, 57] assume the on-shell form for the photon-proton interaction and perform a straightforward evaluation of the loop diagram. Such an evaluation is ultraviolet-finite and collinear-safe, but it violates unitarity for the real parts of the two-photon exchange amplitudes. This violation is not significant for the proton, or elastic, intermediate state, when the lepton mass is neglected [58, 59], but it leads to cross sections that diverge with beam energy when modeling the resonance intermediate states [60, 61]. The dispersion-relation approach [62, 63, 64, 65, 58, 66, 59], based on the on-shell information only, does not suffer from such pathological behavior and is considered a physically-motivated and favorable framework for the evaluation of the two-photon exchange diagram in electron-proton scattering community [67]. As of today, proton and pion-nucleon intermediate states in the two-photon exchange diagram are included in the dispersion-relation approach. Such a calculation takes experimentally measured proton form factors [4, 5] and invariant amplitudes of pion electroproduction [68, 69] as an input. Accounting for the pion-nucleon intermediate states improves agreement between theoretical predictions and recent experimental data from electron- and positron-proton scattering experiments [70, 71, 72, 73, 74, 75, 76]. The remaining discrepancy might be explained by the resonant enhancement of nucleon excitations at GeV beam energies [77], when the propagator of the resonance intermediate state moves close to its pole position.
@@
A similar diagram with the exchange of two bosons contributes to the elastic (anti)neutrino-nucleon scattering. To fully benefit from experimental advances in the fields of neutrino scattering and oscillation experiments [78, 79, 80], this diagram was recently revisited by a few independent groups [81, 82, 83, 84, 85]. As a result, it was included in a new and complete factorization framework for radiative corrections to the charged-current elastic (anti)neutrino-nucleon scattering in Refs. [84, 85]. The latter references evaluated the nucleon intermediate state and provided a model consistent with the expected soft and collinear behavior. The authors estimated the uncertainty from inelastic intermediate states as a contribution from the diagrams with electromagnetic coupling to the neutron. As in electron scattering, resonance intermediate states in the charged-current (anti)neutrino elastic scattering might be enhanced for the incoming beam energies that move the resonance propagators close to their pole positions. That is why resonance contributions may be relevant for precision analyses of (anti)neutrino scattering data.
@@
In this paper, we perform the first evaluation of the $\Delta(1232)$-resonance contribution to the unpolarized charged-current elastic (anti)neutrino-nucleon scattering at GeV energies. We verify that fixed-$t$ dispersion relations do not apply to resonance contributions as they are for the elastic intermediate state [84, 85]. Therefore, we calculate the loop integrals by modeling the $\Delta$ interaction vertex in its on-shell form with two choices: (1) shifting momenta to satisfy collinear constraints as in the model of Refs. [84, 85] and (2) retaining the full momentum dependence in the loop integration as in the hadronic model of Refs. [29, 51, 52, 53, 54, 55, 56, 57], cf. also an interesting discussion in Ref. [86]. Our model satisfies soft and collinear constraints on the real parts of the invariant amplitudes. It also respects the expected crossing properties for several contributions, specified in Section 4, and fulfills the unitarity constraints for the imaginary parts. Additionally, we investigate soft and collinear properties of the invariant amplitudes and study off-shell effects by changing the vector part of the weak interaction vertex of the $\Delta$ resonance from the gauge-independent form of the magnetic dipole contribution only to a traditional in neutrino physics form-factor decomposition.
@@
The paper is organized as follows. In Section 2, we describe the production of the $\Delta$ resonance in electroweak processes at tree level and specify the amplitudes for the $\Delta$ neutrinoproduction. We provide the invariant amplitude decomposition and general cross-section expression in the elastic (anti)neutrino-nucleon scattering in Section 3. In Section 4, we evaluate the virtual contribution of the $\Delta$ intermediate state to the elastic (anti)neutrino-nucleon scattering amplitudes and cross sections and provide numerical results for incoming (anti)neutrino beam energies around GeV scale. In Section 5, we discuss $\Delta$ intermediate state contribution to radiative corrections in the factorization framework and corresponding radiation of real photons. Conclusions and outlook are presented in Section 6. In Appendix A, we specify relations between invariant and helicity amplitudes in the elastic (anti)neutrino-nucleon scattering.
@@ 2 $\Delta$ production in charged-current (anti)neutrino-nucleon scattering
We consider all possible channels for the production of the $\Delta$ resonance in the charged-current scattering of the neutrino $\nu_{\ell}$ and antineutrino $\overline{\nu}_{\ell}$ on the neutron $n$ and proton $p$: 1) $\nu_{\ell}n\to\ell^{-}\Delta^{+}$; 2) $\nu_{\ell}p\to\ell^{-}\Delta^{++}$; 3) $\overline{\nu}_{\ell}n\to\ell^{+}\Delta^{-}$; 4) $\overline{\nu}_{\ell}p\to\ell^{+}\Delta^{0}$. We denote the momentum of the incoming (anti)neutrino as $p$, the momentum of the outgoing charged lepton as $p^{\prime}$, the momentum of the initial-state nucleon as $k$, and the momentum of the final-state resonance as $k^{\prime}$. The kinematics of the resonance production can be completely described by the invariant mass of the resonance $W$, where $W^{2}=\left(k^{\prime}\right)^{2}$, and two Mandelstam variables: the squared momentum transfer $Q^{2}=-\left(k^{\prime}-k\right)^{2}$ and the squared energy in the (anti)neutrino-nucleon center-of-mass reference frame $s=\left(k+p\right)^{2}$.
@@
where the neutron electric form factor is taken from Ref. [110]. For the neutron magnetic, proton electric, and proton magnetic form factors, a dipole form is assumed. Using the magnetic dipole transition approximation of Eq. (5), the coupling to virtual photons becomes gauge-invariant and the unpolarized $\Delta$ production cross section does not contain vector-axial-vector interference terms. In this approximation, the vector transition form factors in Eq. (3) are given by
@@
The conservation of the vector current implies $C^{V}_{6}=0$.
@@
which we call the $\Delta_{\mathrm{I}}$ model. We also consider an alternative approximation $C^{A}_{4}\left(Q^{2}\right)=0$ of Ref. [113], which we call the $\Delta_{\mathrm{II}}$ model. For numerical estimates in this paper, we take the axial-vector transition form factor $C^{A}_{5}\left(Q^{2}\right)$ from Ref. [106]:
@@
3 Charged-current elastic (anti)neutrino-nucleon scattering
In this Section, we describe the invariant amplitude decomposition in the elastic (anti)neutrino-nucleon scattering for the charged-current processes.
@@ 3 Charged-current elastic (anti)neutrino-nucleon scattering
with the averaged nucleon momentum $K_{\mu}=(k_{\mu}+k^{\prime}_{\mu})/2$. Accounting for the nonzero lepton mass, four more invariant amplitudes are required to describe the matrix elements of the charged-current elastic processes, i.e., $T_{\nu_{\ell}n\to\ell^{-}p}=T^{m_{\ell}=0}_{\nu_{\ell}n\to\ell^{-}p}+T^{m_{\ell}\neq 0}_{\nu_{\ell}n\to\ell^{-}p}$ and $T_{\overline{\nu}_{\ell}p\to\ell^{+}n}=T^{m_{\ell}=0}_{\overline{\nu}_{\ell}p\to\ell^{+}n}+T^{m_{\ell}\neq 0}_{\overline{\nu}_{\ell}p\to\ell^{+}n}$ [118, 119]:
@@
4 $\Delta$ intermediate state in charged-current elastic (anti)neutrino-nucleon scattering
@@
Figure 1: Contributions of the $\Delta$ resonance to the charged-current elastic (anti)neutrino-nucleon scattering are shown. The photon is exchanged between the charged lepton and $N\to\Delta$ transition vertex.
@@
$\mathfrak{Re}\left[f^{\mathrm{x}}_{A},f^{\mathrm{x}}_{P},f^{\mathrm{x}}_{R},f^{\mathrm{x}}_{3}\right]\left(\nu,\penalty 10000\ t\right)$ $=\mathfrak{Re}\left[f^{\mathrm{d}}_{A},f^{\mathrm{d}}_{P},f^{\mathrm{d}}_{R},f^{\mathrm{d}}_{3}\right]\left(-\nu,\penalty 10000\ t\right).$ (30)
@@
The results for $\Delta^{0}_{\mathrm{I}}$ and $\Delta^{0}_{\mathrm{II}}$, calculated using different axial-vector transition form factors, cf. Eq. (12) and the description above, are in very close agreement. This indicates the dominant role of the electroweak axial-vector contribution from the form factor $C_{5}^{A}$. The subtraction of power-suppressed collinear contributions has a negligible impact on the results, which validates the approximate flavor independence for the resonance contributions to the unpolarized cross section. For the electron flavor, this subtraction is negligible compared to the magnitude of the $\Delta$-resonance contribution. However, it is more significant for the muon flavor due to power corrections that involve the muon mass. To explicitly study the mass dependence, we subtract the power-suppressed collinear region from each loop integral in the default model while keeping the kinematic relations between the amplitudes and loop integrals exact. We refer to this calculation as $\Delta^{0}_{\mathrm{I}},\penalty 10000\ \mathrm{no\penalty 10000\ coll.}$”, and compare it with the results obtained without any subtraction.
Nevertheless, the $\Delta(1232)$-resonance contribution demonstrates an expected kinematic enhancement, as shown in Figures 2-5, the resulting correction does not exceed the permille level. For the dominant region of (anti)neutrino fluxes in oscillation experiments, the kinematic enhancement happens away from the near-forward region, $Q^{2}\approx 0$, with the largest cross sections. For further validation, we also perform a calculation in the hadronic model $\Delta^{\mathrm{hm}}_{\mathrm{I},\mathrm{II}}$ and find a permille-level correction with a larger kinematic enhancement in a relatively limited phase-space region. The differences between model calculations for the $\Delta$-resonance contribution to the unpolarized elastic (anti)neutrino-nucleon scattering are consistent with the uncertainty estimates in Refs. [84, 85].
@@
Figure 2: Relative contribution of the $\Delta$ intermediate state to the unpolarized charged-current muon neutrino-neutron elastic scattering cross section is shown as a function of the squared momentum transfer $Q^{2}$ for initial neutrino energies $E_{\nu}=600\penalty 10000\ \mathrm{MeV}$ (upper left plot), $E_{\nu}=1\penalty 10000\ \mathrm{GeV}$ (upper right plot), $E_{\nu}=1.2\penalty 10000\ \mathrm{GeV}$ (lower left plot), and $E_{\nu}=2\penalty 10000\ \mathrm{GeV}$ (lower right plot). Results for different model calculations, which are described in Section 4, are presented.
Figure 3: Same as Figure 2 but for the antineutrino-proton scattering.
@@
Our study without enforcing the consistent couplings [122, 123, 108, 109] finds a 11-33 permille-level dependence on the treatment of off-shell contributions in the radiative $\Delta$ decay, which rules out predictions based on the gauge-dependent decomposition of vector contributions in Eq. (3) more precise than order-of-magnitude estimates for both moderate and large momentum transfers. This result is in qualitative agreement with findings for the radiative $\Delta$ decay in neutral-current process [124, 125]. We therefore provide an independent uncertainty estimate for background channels in searches for new physics at short-baseline neutrino experiments [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150].
5 Resonance contributions to radiative corrections in factorization framework
This Section presents the relationship between the present calculation and the factorization approach to QED radiative corrections for charged-current elastic (anti)neutrino-nucleon scattering developed in Refs. [84, 85].
In this framework, applicable at accelerator (anti)neutrino energies, the differential cross section $\mathrm{d}\sigma$ factorizes into hard ($H$), jet ($J$), and soft ($S$) functions as
$\mathrm{d}\sigma\sim H\left(\frac{\mu}{\Lambda}\right)J\left(\frac{\mu}{m_{\ell}}\right)S\left(\frac{\mu}{\Delta E}\right),$ (33)
where $\mu$ is the common renormalization scale, $\Lambda\sim M_{\Delta}\sim M\sim E_{\nu}\sim\sqrt{Q^{2}}$ is the hard scale, and $\Delta E$ is the soft-photon energy acceptance parameter, which defines the maximum energy of photons treated as soft radiation. The soft and jet functions are perturbative and encapsulate the universal long-distance physics responsible for large logarithmic enhancements, which correspond to infrared and collinear singularities. These logarithms become significant at the hard scale $\mu\sim\Lambda$. In contrast, the hard function is nonperturbative and describes QED interactions of hadrons.
At leading order, the hard function $H^{\mathrm{0}}$ is expressed in terms of nucleon form factors. The dominant QED correction $H^{\mathrm{N}}$ from the nucleon intermediate states was computed in Refs. [84, 85]. The primary goal of this work is to calculate the contribution of the $\Delta(1232)$ resonance, $H^{\Delta}$, thereby extending the hard function as
$H=H^{\mathrm{0}}+H^{\mathrm{N}}+H^{\mathrm{\Delta}}.$ (34)
Our analysis demonstrates that the magnitude of $H^{\mathrm{\Delta}}$ is comparable to the dominant theoretical uncertainty associated with inelastic intermediate states in the previously calculated $H^{\mathrm{0}}+H^{\mathrm{N}}$ in Refs. [84, 85]. The logarithmic enhancement from $H^{\mathrm{\Delta}}$ enters the scattering cross section as power-suppressed contributions, resulting in a permille-level virtual correction. An analogous suppression pattern also applies for the $\Delta$-induced radiation of real photons, as these contributions are free from soft and collinear singularities at $\mathcal{O}\left(e^{2}\right)$. Consequently, while the $\Delta$ resonance constitutes a non-negligible component of the hard function, its impact on the overall QED radiative corrections remains sub-dominant compared to the leading logarithmic enhancements captured by the functions $S$ and $J$.